OpenStudy (anonymous):

What is the density of NH3 at 800 torr and 25 degrees Celsius?

OpenStudy (ivancsc1996):

Is it a gas?

OpenStudy (anonymous):

Were supposed to use Ideal Gas Law

OpenStudy (anonymous):

@ivancsc1996

OpenStudy (ivancsc1996):

This are the maths, if you have a question please tell me.\[PV=nRT;n=\frac{ m }{ P _{m} }\rightarrow PV=\frac{ m }{ P _{m} }RT;\rho=\frac{ m }{ V }\rightarrow PP _{m}=\rho RT\]Where m=mass, Pm=molar weight, R=Universal gas constant, P=Presion, V=volume, rho=density, n=moles Then you solve for rho:\[\rho=\frac{ PP _{m} }{ RT }=\frac{ 800torr \frac{ 1.013\times10^{5}Pa }{ 760torr }17.03\frac{ g }{ mol }\frac{ 1kg }{ 1000g } }{ 8.314\frac{ J }{ molK } (25+213.15)K}=0.7326\frac{ kg }{ m ^{3} }=\]